# Expanded Generalized Needlet Internal Linear Combination (eGNILC) Framework for the 21-cm Foreground Removal

### v1

@@ 4 Demonstration
Figure 2: Demonstrations of the RPCA and foreground effects on the GNILC foreground removal method with $20$ channels. The left panel shows the 21-cm angular power spectrum at $1252.5\,\mathrm{MHz}$, where the blue dashed line is the input power spectrum, the orange (green) solid line is the power spectrum recovered from “Extrapolation” power-law foregrounds without (with) RPCA. The red (purple) solid line is recovered from “Extrapolation” and “Unresolved” foregrounds combined without (with) RPCA. The dot-dashed lines in the lower region with the same color are the corresponding fractional differences of power spectra ($\Delta C_{\ell}/C^{\rm real}_{\ell}$) with respect to the input. The right panel is a pixel-by-pixel comparison ($\sim 100,000$ pixels) between the input map and the recovered map with the “Extrapolation” foreground model but without the RPCA method (its power spectrum is shown in the orange line on left panel).
@@
Figure 4: Fluctuations of the recovered power spectrum by the plain GNILC in $1000$ realizations of Hi  21-cm intensity maps, with the same foreground model and mask as in Fig. 1. The simulation runs with $20$ frequency channels and contains power-law and unresolved foregrounds. The cosmic variance (grey band) is calculated as $\Delta C_{\ell}/C_{\ell}=\sqrt{2/[f_{\mathrm{sky}}(2\ell+1)]}$, where $f_{\mathrm{sky}}=0.5$ is the sky coverage factor.
We simulate foregrounds and Hi  signals in the frequency range $960\,\mathrm{MHz}$-$1260\,\mathrm{MHz}$ with $20$ channel bins evenly spaced. We switch the foreground between “Extrapolation” and “Extrapolation + Unresolved”, the method between “RPCA” and “Simulation” to examine the foreground cleaning effect. Figure 2 shows the recovered 21-cm auto-power spectra centered at $1252.5\,\mathrm{MHz}$ in four cases. “Extrapolation” denotes the well-modelled sources in Sec. 2.2.1, and “Unresolved” denotes the unresolved random sources shown in Sec. 2.2.2. The GNILC method and the RPCA-embedded GNILC method are utilized in the foreground removal, which are denoted by “w.o. RPCA” and “w. RPCA” separately according to the manner in which the covariance matrix is computed. Because almost all 21-cm IM experiments observe the partial sky, in the following comparison, we just compare the pseudo-power spectrum (L.H.S. of Eq. (9)) directly, without inverting the $M$-matrix (Eq. (10)) in Eq. (9) to recover the full-sky power spectrum. In a word, we will suppress the tilde in $\tilde{C}_{\ell}$ in the following presentation.
Figure 2’s left panel shows the 21-cm angular power spectrum at $1252.5\,\mathrm{MHz}$, where the blue dashed line is the input power spectrum, and the orange and green solid lines are the power spectrum recovered from “Extrapolation” power-law foregrounds without and with RPCA respectively. The red solid line is the recovered power spectrum from “Extrapolation” and “Unresolved” foregrounds without combining with RPCA, and the purple solid line is that combined with RPCA. In the lower region of the left panel, the dot-dashed lines in the same color are the corresponding fractional differences of power spectra with respect to the input power spectrum. For clarity of presentation, all power spectra curves are smoothed properly to remove the small wiggles in $\ell$-space. The right panel makes a pixel-by-pixel comparison between the recovered Hi  map and the input map, in which the recovered map is the case of “Extrapolation” power-law foregrounds without RPCA. The grey dashed line is the $y=x$ diagonal line, which indicates the unbiased estimate. One can see that the recovered and input values are highly aligned alongside of the $y=x$ line, which demonstrates the level of unbiasedness in the real space (the left panel is the Fourier space).
@@
We also studied the stability of the GNILC method. Because the blind method is comparable to the plain GNILC method regarding accuracy, we only estimated the fluctuations of the plain GNILC method to save computing time. The sky map is generated with fixed foreground maps, and the 1000 Hi  map samples are generated from the same power spectrum. Then the Hi  signal is extracted with the GNILC method from each sample. In Fig. 4, the blue curve is the theoretical angular power spectrum of Hi  (input), the yellow curve is the mean value of the recovered angular power spectrum of Hi , and the green curve and band are the difference and standard deviation. We calculated the cosmic variance $\Delta C_{\ell}/C_{\ell}=\sqrt{2/[f_{\mathrm{sky}}(2\ell+1)]}$ (grey shadow, $f_{\mathrm{sky}}=0.5$ is the sky coverage factor) and put the centre value on the green curve for a vision comparison. The uncertainty caused by the GNILC method is insignificant because the fluctuation of the power spectrum does not override the cosmic variance.
@@ 5 Application to single-dish experiments
We implement the eGNILC foreground removal framework in the simulated data for two single-dish experiments, the SKA-MID and BINGO, both of which probe the low-$z$ 21-cm power spectrum. The DCT transform and the modified AIC is used, and the RPCA is switched off to save computation time. We limit the sky coverage in accordance with the experimental designs and take into account the beam effect, with experimental specifications shown in Table 1. The Airy beam is truncated at $6$ times the radius of the first zero, and the radius is set by best-fitting the parameter $\theta_{\text{FWHM}}$. With this truncation, the last sidelobe is $40\,\mathrm{dB}$ below the maximum, but one should notice that the real primary beam could exhibit much more complex patterns, such as those investigated in Asad et al. (2021) for the MeerKAT (SKA-MID precursor) L-band beams. The Airy power pattern is (Condon & Ransom, 2016)
@@
where $J_{1}$ is the first order Bessel function of the first kind, $u=\sin\theta$ is the sine of the zenith angle, $\lambda$ is the wavelength, and $D$ is the aperture width. Therefore, the beam size is scaled by the frequency ratio $\nu_{0}/\nu$ where $\nu_{0}$ is a reference frequency, which may cause significant errors on typical scales in the recovery. We will see abrupt peaks in the recovered angular power spectra of the SKA-MID and BINGO in Fig. 6 and Fig. 7. The overall system temperature consists of the receiver noise temperature, the temperature of the CMB, the galaxy temperature, etc. As we stated in Sec. 2.2.3, the noise effect is not the major issue of this paper, and we will only extract the 21-cm signal from the foreground contaminants. The potential RFI contamination is also ignored.
To recover the signal with good accuracy, we should deconvolve the native beams and reconvolve a common beam to all maps (Delabrouille et al., 2009). However, in this section, we do not deconvolve the beam from the simulated data before doing the foreground removal. Perfect beam deconvolution is not easy due to non-negligible instrumental noise, scanning strategy and map making, and beam distortion (Burigana & Saez, 2003; Ade et al., 2016). In the real data, it is possible that the beam effect persists on small scales after beam deconvolution. We will present how the eGNILC works in a bad situation where frequency-dependent Airy beams are convolved to maps and beam deconvolution is absent, although the instrument noise is not added.
Figure 5: The sky coverage of the simulated SKA-MID and BINGO, projected in the Equatorial coordinate. The grey area is masked, which is derived from Haslam map. The valid area is reduced from the survey area of $20000\deg^{2}$ to $10000\deg^{2}$ and from the survey area of $5000\deg^{2}$ to $3000\deg^{2}$ (Table 1), respectively, because of the Galactic plane. See also Sec. 5.1.
@@ 5.1 SKA-MID
SKA-MID is a dish array consisting of a set of subarrays, which comprises the South African SKA precursor MeerKAT (launched in 2018) with 64 $13.5\mathrm{m}$-diameter dishes and the supplement of $133$ $15\mathrm{m}$-diameter SKA1 dishes (Bacon et al., 2018). The telescopes are configured with a compact core and three logarithmically spaced spiral arms with a maximum baseline of $150$ km, which corresponds to an angular resolution $0.3\,\mathrm{arcsecs}$ at the frequency of $1.4\,\mathrm{GHz}$ (Bacon et al., 2018). However, SKA-MID in interferometric mode does not provide enough short baselines to map the scales of interest with sufficient signal-to-noise (Bull et al., 2015). We use the single-dish mode instead, which generates the auto-correlation information from each dish [^5] The single-dish IM is an auto-correlation of each dish’s voltage, which is different from the interferometric mode where the visibility is obtained by cross-multiplying the voltage.. The large number of dishes available with the telescopes will guarantee a large survey speed for probing the Hi  signal (Santos et al., 2015).
@@
We apply the eGNILC algorithm to the simulated SKA-MID data. The foreground consists of the “Extrapolation” with or without the “Unresolved” components. Because the frequency resolution is $\delta\nu=0.2\,$MHz (Table 1), the neighbouring 21-cm maps are highly correlated [^6] One may refer to Shaw & Lewis (2008) and Li & Ma (2017) for detailed calculations of the neighbouring frequency band cross-correlation.. This may potentially cause practical difficulties while inverting the signal covariance because of the small eigenvalues of $\bm{R}_{\bm{s}}$. If so, then the computational error is out of control, and consequently, the GNILC foreground removal fails in its algorithm. The solution is to separate the maps into $n$ subsets. For the full set indexed with $\{1_{1},1_{2},\ldots,1_{n};2_{1},2_{2},\ldots,2_{n};\cdots;i_{1},i_{2},\ldots,i_{n};\cdots\}$, $n$ should be carefully assigned to guarantee that all maps in the subset $\{1_{k},2_{k},\cdots,i_{k},\cdots\}$ are uncorrelated. We set $n$ to $10$ for the $500$ frequency channels, which is adequate to break the correlation among 21-cm maps within the subsets and is large enough to obtain high recovery accuracy. The eGNILC is applied to each subset to extract the HI signal.
@@
Figure 6 presents the recovered HI 21-cm angular power spectra at $1069.9\,\mathrm{MHz}$ and $1069.7\,\mathrm{MHz}$ that belong to two different subsets. The original 21-cm maps, together with the foreground, are then convolved with the Airy-disk primary beam to mimic the measurement, and the recovered HI 21-cm maps are compared with the input maps which are the masked convolved 21-cm maps. For all maps, we multiplied the same reprocessed mask from Haslam $408\,\mathrm{MHz}$ map (Remazeilles et al., 2015). In all three panels of Fig. 6, the black thin curves are their original power spectra without the beam, and the blue ones are the ones convolved with the beam (input) so they are damped at large $\ell$s. The orange thick curves are the recovered power spectra with the spectral window function parameter $B=2.5$, while the green thin curves are recovered with $B=1.7$. The pink thick curves and the olive thin curves show their relative differences between the input and recovered power spectrum, whose values can be read out from the right boundary line. In addition, the 21-cm power spectra recovered from the “Extrapolation” and “Extrapolation + Unresolved” foregrounds are plotted in solid and dashed curves respectively.
@@
From the lower left panel of Fig. 6, one can see that there is a strong correlation between the 21-cm maps from two adjacent frequency channels. The difference between “Input” and “Recovery” is $10\%\sim 20\%$ (pink thick lines and olive thin lines; also see Table 2), which shows the good agreement between the two signals. Therefore, the eGNILC can recover the 21-cm signal very well within $\ell=[20,250]$. The truncated Airy beam decreases the power spectrum exponentially with a flat tail. For $\ell\gtrsim 250$, the recovered 21-cm power spectra deviate from the original power spectra because the beam varies with frequency, where larger errors are observed in the recovered power spectrum. Such a region is marked with a grey shadow and is not used in further analysis. Similarly, the left-most region $\ell\lesssim 20$ is also marked with a grey shadow for its large bias due to insufficient independent samples for the eGNILC. We change the parameter $B$ and foreground contaminants to see how the recovered 21-cm is affected. Figure 6 shows the power spectrum bias, and Table 2 gives the average absolute difference in $\ell=[20,250]$ between the input power spectrum and the recovered power spectrum. By changing the foregrounds from the “Extrapol” to “Extrapol + Unresol”, we see the bias increases with the foreground complexity, though the difference is tiny in the auto-power spectra at $1069.9\mathrm{MHz}$. Changing $B$ from $2.5$ to $1.7$ also improves the performance of the eGNILC.
@@ 5.2 BINGO
BINGO has a better spatial resolution than SKA-MID because of its larger dish. In this paper, we simulate its performance in the frequency range of $960\,\mathrm{MHz}$-$1260\,\mathrm{MHz}$ with channel width $\delta\nu=5\,\mathrm{MHz}$, which corresponds to the redshift range of $z=0.13$-$0.48$. The sky coverage of the simulated BINGO data is about $3000\deg^{2}$, and Figure 5 shows its valid area. We take the values of these telescope parameters as per guidelines in the previous literatures (Battye et al. 2012; Dickinson 2014; Battye et al. 2016; Bacon et al. 2018; Wuensche 2019; Yohana et al. 2019; Wang et al. 2021; Abdalla et al. 2022; Wuensche et al. 2022; see also Table 1), but it just serves as a reference to present the characteristics of the two telescopes. The final telescope can have different experimental specifications than the values quoted here. Figure 7 presents the results of foreground removal and 21-cm signal recovery for BINGO in the same way as Figure 6. Because BINGO’s frequency resolution is $\delta\nu=5$ MHz, the neighbouring 21-cm maps are almost uncorrelated, which shows in the lower-left panel that the cross-correlation is an order of magnitude lower than the auto-correlation (upper and right panels of Fig. 7). The cross-power spectrum is therefore plotted linearly and the relative difference is not presented because of the smallness of the cross powers. Because the resolution of BINGO is higher, the cutoff range due to discordant beams at high $\ell$ shrinks to $\ell\gtrsim 300$. After $B$ is changed from $2.5$ to $1.7$, the average absolute difference and the abrupt peaks decrease though the peaks are still distinct (see Table 2 and Fig. 7). Unexpectedly, the 21-cm signal recovered from the “Extrapol + Unresol” has the smallest bias. This indicates that the performance of the eGNILC rest on many key factors, e.g., mask, beam, foreground, spectral window function, and covariance calculation. Given an observation, we may vary the tuning parameters of the eGNILC to obtain the best performance.

### v2

@@ 4 Demonstration
Figure 2: Demonstrations of the RPCA and foreground effects on the GNILC foreground removal method with $20$ channels. The left panel shows the 21-cm angular power spectrum at $1252.5\,\mathrm{MHz}$, where the blue dashed line is the input power spectrum, the orange (green) solid line is the power spectrum recovered from “Extrapolation” power-law foregrounds without (with) RPCA. The red (purple) solid line is recovered from “Extrapolation” and “Unresolved” foregrounds combined without (with) RPCA. The dotted-dashed lines in the lower region with the same colors are the corresponding fractional differences of power spectra ($\Delta C_{\ell}/C^{\rm real}_{\ell}$) with respect to the input. The right panel shows a pixel-by-pixel comparison ($\sim 100,000$ pixels) between the input map and the recovered map with the “Extrapolation” foreground model but without the RPCA method (its power spectrum is shown in the orange line on left panel).
@@
Figure 4: Fluctuations of the recovered power spectrum by the plain GNILC in $1000$ realizations of Hi  21-cm intensity maps, with the same foreground model and mask as in Fig. 1. The simulation runs with $20$ frequency channels and contains power-law and unresolved foregrounds. The cosmic variance (gray band) is calculated as $\Delta C_{\ell}/C_{\ell}=\sqrt{2/[f_{\mathrm{sky}}(2\ell+1)]}$, where $f_{\mathrm{sky}}=0.5$ is the sky coverage factor.
We simulate foregrounds and Hi  signals in the frequency range $960\,\mathrm{MHz}$-$1260\,\mathrm{MHz}$ with $20$ channel bins evenly spaced. We switch the foreground between “Extrapolation” and “Extrapolation + Unresolved”, the method between “RPCA” and “Simulation” to examine the foreground cleaning effect. Figure 2 shows the recovered 21-cm auto-power spectra centered at $1252.5\,\mathrm{MHz}$ in four cases. “Extrapolation” denotes the well-modeled sources in Sec. 2.2.1, and “Unresolved” denotes the unresolved random sources shown in Sec. 2.2.2. The GNILC method and the RPCA-embedded GNILC method are utilized in the foreground removal, which are denoted by “w.o. RPCA” and “w. RPCA” separately according to the manner in which the covariance matrix is computed. Because almost all 21-cm IM experiments observe the partial sky, in the following comparison, we just compare the pseudo-power spectrum (L.H.S. of Eq. (9)) directly, without inverting the $M$-matrix (Eq. (10)) in Eq. (9) to recover the full-sky power spectrum. In a word, we will suppress the tilde in $\tilde{C}_{\ell}$ in the following presentation.
Figure 2’s left panel shows the 21-cm angular power spectrum at $1252.5\,\mathrm{MHz}$, where the blue dashed line is the input power spectrum, and the orange and green solid lines are the power spectrum recovered from “Extrapolation” power-law foregrounds without and with RPCA respectively. The red solid line is the recovered power spectrum from “Extrapolation” and “Unresolved” foregrounds without combining with RPCA, and the purple solid line is that combined with RPCA. In the lower region of the left panel, the dotted-dashed lines in the same color are the corresponding fractional differences of power spectra with respect to the input power spectrum. For clarity of presentation, all power spectra curves are smoothed properly to remove the small wiggles in $\ell$-space. The right panel makes a pixel-by-pixel comparison between the recovered Hi  map and the input map, in which the recovered map is the case of “Extrapolation” power-law foregrounds without RPCA. The gray dashed line is the $y=x$ diagonal line, which indicates the unbiased estimate. One can see that the recovered and input values are highly aligned alongside of the $y=x$ line, which demonstrates the level of unbiasedness in the real space (the left panel is the Fourier space).
@@
We also studied the stability of the GNILC method. Because the blind method is comparable to the plain GNILC method regarding accuracy, we only estimated the fluctuations of the plain GNILC method to save computing time. The sky map is generated with fixed foreground maps, and the 1000 Hi  map samples are generated from the same power spectrum. Then the Hi  signal is extracted with the GNILC method from each sample. In Fig. 4, the blue curve is the theoretical angular power spectrum of Hi  (input), the yellow curve is the mean value of the recovered angular power spectrum of Hi , and the green curve and band are the difference and standard deviation. We calculated the cosmic variance $\Delta C_{\ell}/C_{\ell}=\sqrt{2/[f_{\mathrm{sky}}(2\ell+1)]}$ (gray shadow, $f_{\mathrm{sky}}=0.5$ is the sky coverage factor) and put the center value on the green curve for a vision comparison. The uncertainty caused by the GNILC method is insignificant because the fluctuation of the power spectrum does not override the cosmic variance.
@@ 5 Application to single-dish experiments
We implement the eGNILC foreground removal framework in the simulated data for two SD experiments, the SKA-MID and BINGO, both of which probe the low-$z$ 21-cm power spectrum. The DCT transform and the modified AIC is used, and the RPCA is switched off to save computation time. We limit the sky coverage in accordance with the experimental designs and take into account the beam effect, with experimental specifications shown in Table 1. The Airy beam is truncated at $6$ times the radius of the first zero, and the radius is set by best-fitting the parameter $\theta_{\text{FWHM}}$. With this truncation, the last sidelobe is $40\,\mathrm{dB}$ below the maximum, but one should notice that the real primary beam could exhibit much more complex patterns, such as those investigated in Asad et al. (2021) for the MeerKAT (SKA-MID precursor) L-band beams. The Airy power pattern is (Condon & Ransom, 2016)
@@
where $J_{1}$ is the first-order Bessel function of the first kind, $u=\sin\theta$ is the sine of the zenith angle, $\lambda$ is the wavelength, and $D$ is the aperture width. Therefore, the beam size is scaled by the frequency ratio $\nu_{0}/\nu$ where $\nu_{0}$ is a reference frequency, which may cause significant errors on typical scales in the recovery. We will see abrupt peaks in the recovered angular power spectra of the SKA-MID and BINGO in Fig. 6 and Fig. 7. The overall system temperature consists of the receiver noise temperature, the temperature of the CMB, the galaxy temperature, etc. As we stated in Sec. 2.2.3, the noise effect is not the major issue of this paper, and we will only extract the 21-cm signal from the foreground contaminants. The potential RFI contamination is also ignored.
To recover the signal with good accuracy, we should deconvolve the native beams and reconvolve a common beam to all maps (Delabrouille et al., 2009). However, in this Section, we do not deconvolve the beam from the simulated data before doing the foreground removal. Perfect beam deconvolution is not easy due to nonnegligible instrumental noise, scanning strategy and map making, and beam distortion (Burigana & Saez, 2003; Ade et al., 2016). In the real data, it is possible that the beam effect persists on small scales after beam deconvolution. We will present how the eGNILC works in a bad situation where frequency-dependent Airy beams are convolved to maps and beam deconvolution is absent, although the instrument noise is not added.
Figure 5: The sky coverage of the simulated SKA-MID and BINGO, projected in the Equatorial coordinate. The gray area is masked, which is derived from Haslam map. The valid area is reduced from the survey area of $20,000$ to $10,000\deg^{2}$ and from the survey area of $5000$ to $3000\deg^{2}$ (Table 1), respectively, because of the Galactic plane. See also Sec. 5.1.
@@ 5.1 SKA-MID
SKA-MID is a dish array consisting of a set of subarrays, which comprises the South African SKA precursor MeerKAT (launched in 2018) with 64 $13.5\mathrm{m}$-diameter dishes and the supplement of $133$ $15\mathrm{m}$-diameter SKA1 dishes (Square Kilometre Array Cosmology Science Working Group et al., 2020). The telescopes are configured with a compact core and three logarithmically spaced spiral arms with a maximum baseline of $150$ km, which corresponds to an angular resolution $30^{\prime\prime}.$3 at the frequency of $1.4\,\mathrm{GHz}$ (Square Kilometre Array Cosmology Science Working Group et al., 2020). However, SKA-MID in interferometric mode does not provide enough short baselines to map the scales of interest with sufficient signal-to-noise (Bull et al., 2015). We use the single-dish mode instead, which generates the auto-correlation information from each dish [^5] The single-dish IM is an auto-correlation of each dish’s voltage, which is different from the interferometric mode where the visibility is obtained by cross-multiplying the voltage.. The large number of dishes available with the telescopes will guarantee a large survey speed for probing the Hi  signal (Santos et al., 2015).
@@
We apply the eGNILC algorithm to the simulated SKA-MID data. The foreground consists of the “Extrapolation” with or without the “Unresolved” components. Because the frequency resolution is $\delta\nu=0.2\,$MHz (Table 1), the neighboring 21-cm maps are highly correlated [^6] One may refer to Shaw & Lewis (2008) and Li & Ma (2017) for detailed calculations of the neighboring frequency band cross-correlation.. This may potentially cause practical difficulties while inverting the signal covariance because of the small eigenvalues of $\bm{R}_{\bm{s}}$. If so, then the computational error is out of control, and consequently, the GNILC foreground removal fails in its algorithm. The solution is to separate the maps into $n$ subsets. For the full set indexed with $\{1_{1},1_{2},\ldots,1_{n};2_{1},2_{2},\ldots,2_{n};\cdots;i_{1},i_{2},\ldots,i_{n};\cdots\}$, $n$ should be carefully assigned to guarantee that all maps in the subset $\{1_{k},2_{k},\cdots,i_{k},\cdots\}$ are uncorrelated. We set $n$ to $10$ for the $500$ frequency channels, which is adequate to break the correlation among 21-cm maps within the subsets and is large enough to obtain high recovery accuracy. The eGNILC is applied to each subset to extract the HI signal.
@@
Figure 6 presents the recovered HI 21-cm angular power spectra at $1069.9\,\mathrm{MHz}$ and $1069.7\,\mathrm{MHz}$ that belong to two different subsets. The original 21-cm maps, together with the foreground, are then convolved with the Airy-disk primary beam to mimic the measurement, and the recovered HI 21-cm maps are compared with the input maps which are the masked convolved 21-cm maps. For all maps, we multiplied the same reprocessed mask from Haslam $408\,\mathrm{MHz}$ map (Remazeilles et al., 2015). In all three panels of Fig. 6, the black thin curves are their original power spectra without the beam, and the blue ones are those convolved with the beam (input) so they are damped at large $\ell$s. The orange thick curves are the recovered power spectra with the spectral window function parameter $B=2.5$, while the green thin curves are recovered with $B=1.7$. The pink thick curves and the olive thin curves show their relative differences between the input and recovered power spectrum, whose values can be read out from the right boundary line. In addition, the 21-cm power spectra recovered from the “Extrapolation” and “Extrapolation + Unresolved” foregrounds are plotted in solid and dashed curves respectively.
@@
From the lower left panel of Fig. 6, one can see that there is a strong correlation between the 21-cm maps from two adjacent frequency channels. The difference between “Input” and “Recovery” is $10\%\sim 20\%$ (pink thick lines and olive thin lines; also see Table 2), which shows the good agreement between the two signals. Therefore, the eGNILC can recover the 21-cm signal very well within $\ell=[20,250]$. The truncated Airy beam decreases the power spectrum exponentially with a flat tail. For $\ell\gtrsim 250$, the recovered 21-cm power spectra deviate from the original power spectra because the beam varies with frequency, where larger errors are observed in the recovered power spectrum. Such a region is marked with a gray shadow and is not used in further analysis. Similarly, the left-most region $\ell\lesssim 20$ is also marked with a gray shadow for its large bias due to insufficient independent samples for the eGNILC. We change the parameter $B$ and foreground contaminants to see how the recovered 21-cm is affected. Figure 6 shows the power spectrum bias, and Table 2 gives the average absolute difference in $\ell=[20,250]$ between the input power spectrum and the recovered power spectrum. By changing the foregrounds from the “Extrapol” to “Extrapol + Unresol”, we see the bias increases with the foreground complexity, though the difference is tiny in the auto-power spectra at $1069.9\mathrm{MHz}$. Changing $B$ from $2.5$ to $1.7$ also improves the performance of the eGNILC.
@@ 5.2 BINGO
BINGO has a better spatial resolution than SKA-MID because of its larger dish. In this paper, we simulate its performance in the frequency range of $960\,\mathrm{MHz}$-$1260\,\mathrm{MHz}$ with channel width $\delta\nu=5\,\mathrm{MHz}$, which corresponds to the redshift range of $z=0.13$-$0.48$. The sky coverage of the simulated BINGO data is about $3000\deg^{2}$, and Figure 5 shows its valid area. We take the values of these telescope parameters as per guidelines in the previous literatures (Battye et al. 2012; Dickinson 2014; Battye et al. 2016; Square Kilometre Array Cosmology Science Working Group et al. 2020; Wuensche 2019; Yohana et al. 2019; Wang et al. 2021; Abdalla et al. 2022; Wuensche et al. 2022; see also Table 1), but it just serves as a reference to present the characteristics of the two telescopes. The final telescope can have different experimental specifications than the values quoted here. Figure 7 presents the results of foreground removal and 21-cm signal recovery for BINGO in the same way as Figure 6. Because BINGO’s frequency resolution is $\delta\nu=5$ MHz, the neighbouring 21-cm maps are almost uncorrelated, which shows in the lower-left panel that the cross-correlation is an order of magnitude lower than the auto-correlation (upper and right panels of Fig. 7). The cross-power spectrum is therefore plotted linearly and the relative difference is not presented because of the smallness of the cross powers. Because the resolution of BINGO is higher, the cutoff range due to discordant beams at high $\ell$ shrinks to $\ell\gtrsim 300$. After $B$ is changed from $2.5$ to $1.7$, the average absolute difference and the abrupt peaks decrease, although the peaks are still distinct (see Table 2 and Fig. 7). Unexpectedly, the 21-cm signal recovered from the “Extrapol + Unresol” has the smallest bias. This indicates that the performance of the eGNILC rest on many key factors, e.g., mask, beam, foreground, spectral window function, and covariance calculation. Given an observation, we may vary the tuning parameters of the eGNILC to obtain the best performance.